OpenStudy (anonymous):

3. A 0.200 kg air-track glider moving at 1.20 m/s bumps into a 0.600 kg glider at rest. a.) Find the total kinetic energy after collision if the collision is elastic. b.) Find the total kinetic energy after collision if the collision is completely inelastic. Show all your work.

OpenStudy (anonymous):

Mass of the air-track glider ,ma=0.2 kg Speed of the air-track glider ,v1=1.2 m/s Mass of the bumps,mb=0.6 kg Speed of the bumps ,v2=0 a) an elastic collision conserves mechanical energy KE = 1/2 ma* v1^2 Substitute 1.2 for v1 and 0.2 for ma = 1/2 0.200 (1.20)^2 = 0.144 J b) Totally inelastic collisions conservation of momentum M1V1 + M2V2 = (M1+M2)V3 0.200(1.20) + 0.600(0) = 0.800(V3) V3 = 0.300 m/s so KE = 1/2 m v^2 = 1/2 (0.800) (0.300)^2 = 0.036 J